OpenStudy (anonymous):

How would I factor these two polynomials? a) 36x^6-25y^4 b) 4x^6y^4-28x^3y^2+49

OpenStudy (anonymous):

a) \[36x ^{6}-25y ^{4}\] b) \[4x^6y^4-28x^3y^2+49\]

Directrix (directrix):

@rach_ell Try the difference of two squares factoring pattern on this one: 36x^6-25y^4

OpenStudy (anonymous):

I attempted to before, but just wasn't sure what to make of the exponent 6. Usually when it's x^2 or x^4, it's just x or x^2 (respectively) in each bracket .

Directrix (directrix):

How about this: 36x^6-25y^4 = (6x³)² - (5y²)² -> Do you see this as the difference of two squares? @rach_ell

Directrix (directrix):

(6x³)² - (5y²)² =

OpenStudy (anonymous):

Would it be like this? (6x³)² - (5y²)² = (6x^3-5y^2)(6x+5y^2)

Directrix (directrix):

Correct.

Directrix (directrix):

The second problem looks like a perfect square trinomial scenario.

Directrix (directrix):

Reminder attached.

Directrix (directrix):

@rach_ell

OpenStudy (anonymous):

Okay, I got this: \[4x^6y^4-28x^3y^2+49 \] = \[(2x^3y^2)^2-28x^3y^2+49\] = \[(2x^3y^2-7)(2x^3y^2+7)\]

OpenStudy (anonymous):

oops, I mean (2x^3y^2-7)^2

Directrix (directrix):

Do you see the bad sign here? See attachment.

Directrix (directrix):

I agree with this --> (2x^3y^2-7)^2

Directrix (directrix):

@rach_ell

OpenStudy (anonymous):

Yes, I realized the mistake after I checked a second time. Thank you so much for helping!

Directrix (directrix):

You are welcome.